Las Vegas Society of Supernatural Investigaton

 

 

 

 

 

 

 

 

 

 

 

 

Weather Conditions:

Temps in the mid 70's, 10% humidity, Barometric pressure 29.59 and Rising, light winds SSW changing to NNW

Geomagnetic Storms:

very quiet

Solar X-Rays:

class A8 level

Moon Phase:

waning gibbous with 68% of the Moon's visible disk illuminated

Equipment Used:

(All Equipment may not be used at each location but is taken)

Camera Equipment:

 Nikon 35mm, Sony Cyber Shot 50, 35mm Night Owl, Polaroid camera, Samsung digimax a 503

EMF Detectors:

Electrosensor EMF detector,  Cell Sensor EMF detector, ELF Zone Electromagnetic Field Detector,  Mulitdetector II High sensitive Electrosmag-Analyser

Thermometers:

RayTek MiniTemp non-contact thermometer, RayTek Raynger ST non-contact thermometer, RadioShack IR thermometer,

EVP Equipment:

Olympus Digital Voice Recorder DS-320, RCA RP50128 Digital Voice Recorder, Omnidirectional detachable microphone

Video Equipment

Sony Night Shot

Other Equipment:

Copper Dowsing Rods, Motion detectors, Motorola Talkabout T5500 radios, 2 military red lens flashlights, 1 Mag light large regular lens, Rayovac small clip on red lens flashlight, Rayovac head lamp flashlight

 Misc. Items

 Opening and closing prayer, identification badge, assorted pens, assorted batteries, hand sanitizer, first aid kit, snake bite kit, 1 Bag sea salt, 1 pair of scissors, 1 tea ball, sage sticks, torch/flame, holy water and oil, film for cameras

 

BACK

Case History

     The family has heard foot steps, voices, objects fell off the fridge and have seen a shadow figure walking.  One of the children has conversations with whoever is here. They feel that it is 2 females and 1 male in their home.

Date Of Investigation: 06/02/2007
Time:
Investigators on Scene: Tina Carlson and Nancy Riggs
Positive Photos: 3
EVPs: 0
Anomalous EMF readings: 1
Anomalous temperature changes: 0
Video: 0

After the investigation, the members felt that the experiences that they had showed signs of paranormal activity. A cleansing was done and no further activity has been reported to LVSSI.

Click on photo to enlarge then click on  photo to return